Dan has 21 coins, all nickels, dimes, and quarters worth $3.35. If he has 4 more quarters than dimes, how many of each coin does he have?
Answer by TimothyLamb(4379)

x = nickels
y = dimes
z = quarters
---
x + y + z = 21
5x + 10y + 25z = 335
z = y + 4
---
put the system of linear equations into standard form
---
x + y + z = 21
5x + 10y + 25z = 335
-y + z = 4
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = nickels = 5
y = dimes = 6
z = quarters = 10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php